An insulated rod of length has its ends A and B maintained at O^ C and 100 degrees * C respectively until steady state conditions prevailFind the steady -state temperature of the rod


1
Expert's answer
2022-02-15T13:21:32-0500

Let us remind the heat transfer equation :

"\partial_t T = a \partial_{xx}T" as the problem has only one spatial dimension.

The steady state corresponds to the condition "\partial_t T = 0; \; T(t,x)=T(x)". Using this gives us

"\frac{d^2}{dx^2} T = 0"

And the general solution of such equation is given by

"T = C_1x + C_2"

Let us place the origin at the point "A" and let us denote the rod's length by "L". Then we have :

"C_2 = 0^\circ C"

"C_1L+0 = 100 \rightarrow C_1 = 100/L"

Therefore, the temperature is given by "T(x) = \frac{x}{L}\cdot 100"